if Sin^-1x+sin^-1y=2π÷3. Then what is the value of
cos^-1+cos^-1=?

Grade:12th pass

2 Answers

Arun
5 years ago
Dear student
 
cos^-1 (-1) + cos^- 1 (-1) = pi + pi = 2 * pi
 
Hpe it helps
 
Regards
Arun (askIITians forum expert)
Susmita
5 years ago
sin-1x+sin-1y=2pi/3
or,(pi/2-cos-1x)+(pi/2-cos-1y)=2pi/3
Or,cos-1x+cos-1y=pi-2pi/3=pi/3
 
Hope this helps.please approve if it does.